# On gaps in the spectra of quasiperiodic Schrödinger operators with discontinuous monotone potentials

### v2

@@ 3.1. Potentials with infinite values
The standard diagonal procedure argument immediately applies the converse statement.
@@ 4.1. General properties of gap filling
In the natural topology, $\overline{\mathbb{R}}$ is homeomorphic to a circle. We will assume that positive orientation is inherited from that of $\mathbb{R}$. For $t_{1},t_{2}\in\overline{\mathbb{R}}$, a positively-oriented arc that starts at $t_{1}$ and ends at $t_{2}$, will be denoted by $(t_{1},t_{2})$ and called an interval. One can also consider closed intervals $[t_{1},t_{2}]$. For example, one has have
@@
be a spectral gap of $A$, including, possibly, one infinite gap considered as a subset of $\overline{\mathbb{R}}$ (in other words, $(g_{j}^{-},g_{j}^{+})$ are the connected components of the complement of the closure of $\sigma(A)$ in $\overline{\mathbb{R}}$). The general theory of rank one perturbations (see, for example, Chapter 11 of [17]) implies that $\sigma(A_{t})$ with $t\in\overline{\mathbb{R}}\setminus\{0\}$ contains $\sigma(A)$ as well as at most one simple eigenvalue in each $(g_{j}^{-},g_{j}^{+})$. This eigenvalue is described by a continuous monotone function $\lambda_{j}\colon(t_{j}^{-},t_{j}^{+})\to\overline{\mathbb{R}}$, where $(t_{j}^{-},t_{j}^{+})\subset\overline{\mathbb{R}}\setminus\{0\}$ is a (maximal open) interval on which $(g_{j}^{-},g_{j}^{+})$ contains an eigenvalue: that is,
@@ 4.2. Statements of the results
We will consider Schrödinger operators associated to $1$-periodic maps $f\colon\mathbb{R}\to[-\infty,+\infty)$:
@@
where the operator is considered in the sense of (3.1) for infinite values of $f$. Both $[0,1)$ and [,+)[$\infty$,$\infty$)[ - ∞ , + ∞ ) can be identified with circles, making $f|_{[0,1)}$ a circle map. We will consider a class of $f$ such that the corresponding circle maps have one discontinuity (which, without loss of generality, can be placed at the origin).
@@
We say that a $1$-periodic function $f\colon\mathbb{R}\to[-\infty,+\infty)$ has a simple discontinuity if $f|_{[0,1)}$ extends to a continuous map from from $[0,1]$ to [,+][$\infty$,$\infty$][ - ∞ , + ∞ ], with $f(0)\neq f(1-0)$ and at least one value among $f(0)$ and $f(1-0)$ being finite.
Every sawtooth-type potential defined in the introduction has a simple discontinuity. The Maryland-type ones do not, since the corresponding circle maps will be continuous. Let $\mathcal{A}$ be a closed arc in $\overline{\mathbb{R}}$ (which is associated to [,+)[$\infty$,$\infty$)[ - ∞ , + ∞ )) whose endpoints are $f(0)$ and $f(1-0)$. Note that there are two choices of such arc (more on this later). Our goal will be to consider a family of operators, parametrized by $t\in\mathcal{A}$, where the value $f(0)$ will be replaced by $t$. More precisely, for $t\in\mathcal{A}$, define
@@
Our goal is to state a gap filling result of the kind described in Theorem 1.2. In the case of a sawtooth-type potential, say, $f(x)=\{x\}$, there is an obvious choice between two arcs (the points of the arc should complement the range of $f$). In general, since $f$ is not assumed to be monotone, we will need to specify it more carefully. For an arc $\mathcal{A}$ with endpoints $f(0)$ and $f(1-0)$, let
@@
Clearly, if one now identifies $[-1,1)$ and [,+)[$\infty$,$\infty$)[ - ∞ , + ∞ ) with circles, $\tilde{f}$ will become a continuous circle map. We are now ready to state the gap filling result.
@@
Note that, for every $f$ with a simple discontinuity, at least one choice of the arc $\mathcal{A}$ will produce $\tilde{f}$ not homotopic to a constant. As described in the previous section, (4.4) holds for any family of rank one perturbations with $\mathcal{A}$ replaced by $\overline{\mathbb{R}}$ ([17, Theorem 11.8]).
@@
Suppose that $f\colon\mathbb{R}\to[-\infty,+\infty)$ is $1$-periodic, and the circle map associated to $f|_{[0,1)}$ is continuous and not homotopic to a constant. Define $H(x)$ as in (4.1). Then $\sigma(H(x))=\mathbb{R}$.
@@ 4.3. Structure of the hull
In this section, we will construct a family of operator that will extend the original family $\{H(x)\colon x\in[0,1)\}$ and serve the purposes described above. It will be convenient to construct an auxiliary Cantor set whose gaps are labeled by $\{\mathbf{n}\cdot\alpha\}$, $\mathbf{n}\in\mathbb{Z}^{d}$. Since we will be dealing only with aspects of topology, the exact construction of such a set would not matter (but some maps will be needed to be fixed in a way continuous with respect to $\alpha$).
@@
$h(y):=y+3^{-d}\sum_{\mathbf{n}\colon\{\mathbf{n}\cdot\alpha\}\leq y}2^{-|\mathbf{n}|_{1}}.$
@@
The set $\mathcal{C}_{\alpha}$ is a Cantor set if and only if at least one component of $\alpha$ is irrational. In all cases, the gaps or $\mathcal{C}_{\alpha}$ (that is, connected components of $\mathcal{S}\setminus\mathcal{C}_{\alpha}$) are in one-to-one correspondence with the numbers $\{\mathbf{n}\cdot\alpha\}$. Let $\mathcal{G}_{\mathbf{n}}$ be the gap associated to $\mathbf{n}\in\mathbb{Z}$. Note that, if the components of $\alpha$ are rationally independent, different values of $\mathbf{n}$ may correspond to the same value of $\{\mathbf{n}\cdot\alpha\}$ and therefore to the same gap.
@@
where one has $H(x)=H(x+0)$ for all $x$ and $H(x)=H(x-0)$ for $x\in[0,1)\setminus\mathbb{Z}^{d}\cdot\alpha$. For $s,s_{j}\in\{+0,-0\}$, it is easy to see that $H(x_{j}s_{j})\to H(xs)$ if and only if $x_{j}\to xs$ with $x_{j}\neq x$ for $j$ large enough, or $x_{j}s_{j}=xs$ for $j$ large enough. The latter is equivalent to the topology on $\mathcal{C}_{\alpha}$. ∎
@@ 4.4. Rank one perturbations and the extended operator family
As in the previous section, suppose that $f\colon[0,1)\to\overline{\mathbb{R}}$ has a simple discontinuity. Assume that $\mathcal{A}$ is an arc with endpoints $f(0),f(1-0)$. As in the statement of Theorem 4.2, let
@@ 4.6. The case of rational $\alpha$
For rational $\alpha$ (that is, $\alpha$ with all components being rational), the operator $H(\theta)$ becomes a periodic operator. Suppose that, additionally, $\alpha=(p_{1}/q_{1},\ldots,p_{d}/q_{d})$, with $q_{1},\ldots,q_{d}$ pairwise coprime. Let $q=q_{1}\ldots q_{d}$. There exists $\mathbf{n}_{q}\in\mathbb{Z}^{d}$ such that $\mathbf{n}_{q}\cdot\alpha=1/q\,\,\mathrm{mod}\,\,\mathbb{Z}$. The set $\mathcal{C}_{\alpha}$ becomes a union of $q$ arcs, with the set of gap labels being $0,1/q,2/q,\ldots,(q-1)/q$. The operators $H(\theta)$ and $H(\theta+1/q)$ are unitarily equivalent (here, $H(\theta+1/q)$ is defined using the translation by $\mathbf{n}_{q}$). In terms of variable $x$, we also have $H(x)$ unitarily equivalent to $H(x+1/q)$, as well as $H(x-0)$ to $H(x-0-1/q)$. It is also convenient to represent $\mathcal{C}_{\alpha}$ in terms of variable $x$ as a disjoint union
$[0,1/q]\sqcup[1/q,2/q]\sqcup\ldots\sqcup[(q-1)/q,1],$
where the point $1/q$ on the interval $[0,1/q]$ denotes $H(1/q-0)$, and the one on $[1/q,2/q]$ corresponds to $H(1/q+0)$.
@@
Let $u\colon\mathbb{T}\to U(N)$ be a continuous map from the circle to the group of $N\times N$ unitary matrices, and suppose that the map $t\mapsto\det(u(t))$ has a non-zero winding number. Then
@@
Proof of Theorem 4.13. Consider the map $\theta\mapsto U_{h(\theta)}$. In view of the above, it is sufficient to show that the corresponding determinant map has a non-zero winding number. Since the said winding number is a topological invariant, one can replace $\Delta$ by $s\Delta$, where $s$ continuously changes from $1$ to 00 and reduce it to the case of a diagonal operator. In that case, each diagonal entry can be considered as a scalar unitary operator, with the winding number of the determinant equal to that of the map $\tilde{f}$. As a consequence, each diagonal entry will produce an equal non-zero contribution to the total winding number, which completes the proof.

### v3

@@ 3.1. Potentials with infinite values
The standard diagonal procedure argument immediately implies the converse statement.
@@ 4.1. General properties of gap filling
In the natural topology, $\overline{\mathbb{R}}$ is homeomorphic to a circle. We will assume that positive orientation is inherited from that of $\mathbb{R}$. For $t_{1},t_{2}\in\overline{\mathbb{R}}$, a positively-oriented arc that starts at $t_{1}$ and ends at $t_{2}$, will be denoted by $(t_{1},t_{2})$ and called an interval. One can also consider closed intervals $t_{1},t_{2}]$. For example, one has
@@
be a spectral gap of $A$, including, possibly, one infinite gap considered as a subset of $\overline{\mathbb{R}}$ (in other words, $(g_{j}^{-},g_{j}^{+})$ are the connected components of the complement of the closure of $\sigma(A)$ in $\overline{\mathbb{R}}$). The general theory of rank one perturbations (see, for example, Chapter 11 of [20]) implies that $\sigma(A_{t})$ with $t\in\overline{\mathbb{R}}\setminus\{0\}$ consists of $\sigma(A)$ and at most one simple eigenvalue in each $(g_{j}^{-},g_{j}^{+})$. This eigenvalue is described by a continuous monotone function $\lambda_{j}\colon(t_{j}^{-},t_{j}^{+})\to\overline{\mathbb{R}}$, where $(t_{j}^{-},t_{j}^{+})\subset\overline{\mathbb{R}}\setminus\{0\}$ is a (maximal open) interval on which $(g_{j}^{-},g_{j}^{+})$ contains an eigenvalue: that is,
@@ 4.2. Statements of the results
[…]
@@
where the operator is considered in the sense of (3.1) for infinite values of $f$. Both $[0,1)$ and $[-\infty,+\infty)$ can be identified with circles, making $f|_{[0,1)}$ a circle map. We will consider a class of $f$ such that the corresponding circle maps have one discontinuity (which, without loss of generality, can be placed at the origin).
@@
[…]
Every sawtooth-type potential defined in the introduction has a simple discontinuity. The Maryland-type ones do not, since the corresponding circle maps will be continuous. Let $\mathcal{A}$ be a closed arc in $\overline{\mathbb{R}}$ (which is associated to $[-\infty,+\infty)$) whose endpoints are $f(0)$ and f(10)f(1-0). Note that there are two choices of such arc (more on this later). Our goal will be to consider a family of operators, parametrized by $t\in\mathcal{A}$, where the value $f(0)$ will be replaced by $t$. More precisely, for $t\in\mathcal{A}$, define
@@
Our goal is to state a gap filling result of the kind described in Theorem 1.2. In the case of a sawtooth-type potential, say, $f(x)=\{x\}$, there is an obvious choice between two arcs (the points of the arc should complement the range of $f$). In general, since $f$ is not assumed to be monotone, we will need to specify it more carefully. For an arc $\mathcal{A}$ with endpoints $f(0)$ and f(10)f(1-0), let
@@
Clearly, if one now identifies [1,1)[-1,1) and $[-\infty,+\infty)$ with circles, $\tilde{f}$ will become a continuous circle map. We are now ready to state the gap filling result.
@@
Note that, for every $f$ with a simple discontinuity, at least one choice of the arc $\mathcal{A}$ will produce $\tilde{f}$ not homotopic to a constant. As described in the previous section, (4.4) holds for any family of rank one perturbations with $\mathcal{A}$ replaced by $\overline{\mathbb{R}}$ ([20, Theorem 11.8]).
@@
[…]
@@ 4.3. Structure of the hull
In this section, we will construct a family of operators that will extend the original family $\{H(x)\colon x\in[0,1)\}$ and serve the purposes described above. It will be convenient to construct an auxiliary Cantor set whose gaps are labeled by $\{\mathbf{n}\cdot\alpha\}$, $\mathbf{n}\in\mathbb{Z}^{d}$. Since we will be dealing only with aspects of topology, the exact construction of such a set would not matter (but some maps will be needed to be fixed in a way continuous with respect to $\alpha$).
@@
$h(y):=y+3^{-d}\sum_{\mathbf{n}\in\mathbb{Z}^{d}\colon\{\mathbf{n}\cdot\alpha\}\leq y}2^{-|n_{1}|-|n_{2}|-\ldots-|n_{d}|},\quad\text{where}\quad\mathbf{n}=(n_{1},\ldots,n_{d}).$
@@
The set $\mathcal{C}_{\alpha}$ is a Cantor set if and only if at least one component of $\alpha$ is irrational. In all cases, the gaps of $\mathcal{C}_{\alpha}$ (that is, connected components of $\mathcal{S}\setminus\mathcal{C}_{\alpha}$) are in one-to-one correspondence with the numbers $\{\mathbf{n}\cdot\alpha\}$. Let $\mathcal{G}_{\mathbf{n}}$ be the gap associated to $\mathbf{n}\in\mathbb{Z}$. Note that, if the components of $\alpha$ are rationally independent, different values of $\mathbf{n}$ may correspond to the same value of $\{\mathbf{n}\cdot\alpha\}$ and therefore to the same gap.
@@
where one has $H(x)=H(x+0)$ for all $x$ and H(x)=H(x0)H(x)=H(x-0) for $x\in[0,1)\setminus\mathbb{Z}^{d}\cdot\alpha$. For $s,s_{j}\in\{+0,-0\}$, it is easy to see that $H(x_{j}s_{j})\to H(xs)$ if and only if $x_{j}\to xs$ with $x_{j}\neq x$ for $j$ large enough, or $x_{j}s_{j}=xs$ for $j$ large enough. The latter is equivalent to the topology on $\mathcal{C}_{\alpha}$. ∎
@@ 4.4. Rank one perturbations and the extended operator family
As in the previous section, suppose that $f\colon[0,1)\to\overline{\mathbb{R}}$ has a simple discontinuity. Assume that $\mathcal{A}$ is an arc with endpoints f(0),f(10)f(0),f(1-0). As in the statement of Theorem 4.2, let
@@ 4.6. The case of rational $\alpha$
For rational $\alpha$ (that is, $\alpha$ with all components being rational), the operator $H(\theta)$ becomes a periodic operator. Suppose that, additionally, $\alpha=(p_{1}/q_{1},\ldots,p_{d}/q_{d})$, with $q_{1},\ldots,q_{d}$ pairwise coprime. Let $q=q_{1}\ldots q_{d}$. There exists $\mathbf{n}_{q}\in\mathbb{Z}^{d}$ such that $\mathbf{n}_{q}\cdot\alpha=1/q\,\,\mathrm{mod}\,\,\mathbb{Z}$. The set $\mathcal{C}_{\alpha}$ becomes a union of $q$ arcs, with the set of gap labels being $0,1/q,2/q,\ldots,(q-1)/q$. The operators $H(\theta)$ and $H(\theta+1/q)$ are unitarily equivalent (here, $H(\theta+1/q)$ is defined using the translation by $\mathbf{n}_{q}$). In terms of variable $x$, we also have $H(x)$ unitarily equivalent to $H(x+1/q)$, as well as H(x0)H(x-0) to H(x01/q)H(x-0-1/q). It is also convenient to represent $\mathcal{C}_{\alpha}$ in terms of variable $x$ as a disjoint union
$0,1/q]\sqcup[1/q,2/q]\sqcup\ldots\sqcup[(q-1)/q,1],$
where the point $1/q$ on the interval $[0,1/q]$ denotes H(1/q0)H(1/q-0), and the one on $[1/q,2/q]$ corresponds to $H(1/q+0)$.
@@
Let $u\colon\mathbb{T}\to U(N)$ be a continuous map from the circle to the group of $N\times N$ unitary matrices, and suppose that the map $t\mapsto\det(u(t))$, considered as a map from $\mathbb{T}$ to $\mathbb{C}\setminus\{0\}$, has a non-zero winding number around the origin. Then
@@
Proof of Theorem 4.13. Consider the map $\theta\mapsto U_{h(\theta)}$. In view of the above, it is sufficient to show that the corresponding determinant map has a non-zero winding number. Since the said winding number is a topological invariant, one can replace $\Delta$ by $s\Delta$, where $s$ continuously changes from $1$ to 0 and reduce the original statement to the case of a diagonal operator. In that case, each diagonal entry can be considered as a scalar unitary operator, with the winding number of the determinant equal to that of the map $\tilde{f}$. As a consequence, each diagonal entry will produce an equal non-zero contribution to the total winding number, which completes the proof.